What will be the equilibrium level of income and level of consumption in equilibrium if C=30+0.8Y and autonomous investment is 50 crore
1
Expert's answer
2020-11-23T05:42:11-0500

Equilibrium level of income

C = 30+0.8Y

I = 50 crore

Y = C+I

Y = 30+0.8Y+50

Y - 0.8Y = 30 + 50

0.2Y = 80

Y = 400


Equilibrium level of consumption

C = 30 + 0.8Y

C = 30 + 0.8(400)

C = 30 + 320

C = 350